# Rethinking Probabilistic Circuit Parameter Learning

### v1

@@ Appendix C Global Parameter Renormalization of PCs
$\frac{\partial\log{p}_{\boldsymbol{\phi}^{\prime}}^{n}(\boldsymbol{x})}{\partial\log{p}_{\boldsymbol{\phi}^{\prime}}^{c}(\boldsymbol{x})}$ $=\frac{{p}_{\boldsymbol{\phi}^{\prime}}^{c}(\boldsymbol{x})}{{p}_{\boldsymbol{\phi}^{\prime}}^{n}(\boldsymbol{x})}\cdot\frac{\partial{p}_{\boldsymbol{\phi}^{\prime}}^{n}(\boldsymbol{x})}{\partial{p}_{\boldsymbol{\phi}^{\prime}}^{c}(\boldsymbol{x})},$
$=\frac{{p}_{\boldsymbol{\phi}^{\prime}}^{c}(\boldsymbol{x})}{{p}_{\boldsymbol{\phi}^{\prime}}^{n}(\boldsymbol{x})}\cdot\theta^{\prime}_{n,c},$
$=\frac{{p}_{\boldsymbol{\phi}}^{c}(\boldsymbol{x})/Z_{\boldsymbol{\phi}}(c)}{{p}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x})/Z_{\boldsymbol{\phi}}(n)}\cdot\theta^{\prime}_{n,c},$
[…]
$=\frac{{p}_{\boldsymbol{\phi}}^{c}(\boldsymbol{x})}{{p}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x})}\cdot\theta_{n,c},$
$=\frac{{p}_{\boldsymbol{\phi}}^{c}(\boldsymbol{x})}{{p}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x})}\cdot\frac{\partial{p}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x})}{\partial{p}_{\boldsymbol{\phi}}^{c}(\boldsymbol{x})},$
$=\frac{\partial\log{p}_{\boldsymbol{\phi}}^{c}(\boldsymbol{x})}{\partial\log{p}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x})}.$
Appendix D Empirical Evaluations
We train an HCLT [Liu and Van den Broeck, 2021] with hidden size 512 on $16\times 16$ aligned patches from the ImageNet32 dataset [Deng et al., 2009]. That is, we partition every $32\times 32$ image (there are three color channels) into four $16\times 16$ patches and treat these as data samples. There are in total $16\times 16\times 3=768$ categorical variables in the PC.
We apply a lossy YCC transformation proposed by Malvar and Sullivan [2003]. Specifically, given a pixel with RGB values $(R,G,B)$, we first normalize them to the range $[0,1]$ by
@@ Appendix D Empirical Evaluations
where $y$, $co$, and $cg$ are all in the range $[-1,1]$. Finally, we quantize the interval $[-1,1]$ into $256$ bins uniformly and convert $y$, $co$, and $cg$ to their quantized version $Y$, $Co$, and $Cg$, respectively. Note that $Y$, $Co$, and $Cg$ are all categorical variables with $256$ categories.
@@
The full-batch EM implementation follows prior work (e.g., Choi et al. [2021], Peharz et al. [2020]).
Mini-Batch EM.
For notation simplicity, we rewrite Equation 6 as
@@
which makes it more consistent with the baseline mini-batch EM algorithm. For both our algorithm and the baseline algorithm in Equation 5, we choose a base learning rate of $\alpha=0.4$ and a cosine learning rate decay schedule to decrease it to $\alpha=0.08$. For both algorithms, we use a batch size of $32768$. We also use momentum to update the flows $\mathtt{F}_{\boldsymbol{\phi}}^{\mathcal{D}}$. Specifically, we initialize the momentum flows $\mathtt{Fm}_{\boldsymbol{\phi}}^{\mathcal{D}}=\mathbf{0}$, then before every EM step, we update $\mathtt{Fm}_{\boldsymbol{\phi}}^{\mathcal{D}}$ following:
$\mathtt{Fm}_{\boldsymbol{\phi}}^{\mathcal{D}}\leftarrow\eta\cdot\mathtt{Fm}_{\boldsymbol{\phi}}^{\mathcal{D}}+(1-\eta)\cdot\mathtt{F}_{\boldsymbol{\phi}}^{\mathcal{D}}.$
Finally, we replace the $\mathtt{F}_{\boldsymbol{\phi}}^{\mathcal{D}}(n,c)$ in Equation 18 with $\mathtt{Fm}_{\boldsymbol{\phi}}^{\mathcal{D}}(n,c)/(1-\eta^{T+1})$, where $T$ is the number of updates performed.
@@
Following Loconte et al. [2025, 2024], we adopt the Adam optimizer [Kingma, 2014]. We tested with learning rates $21e-$ 2, $33e-$ 3, and $31e-$ 3, and observed that $21e-$ 2 performed the best. This matches the observation in Loconte et al. [2024]. We also tested different batch sizes, and $1024$ performs the best given a fixed number of epochs.
Figure 2: Validation log-likelihood of an HCLT PC [Liu and Van den Broeck, 2021] on $16\times 16$ patches from ImageNet32. Our proposed mini-batch EM algorithms outperform all other optimizers by a large margin.
Empirical Insights.
To have a fair comparison between full-batch and mini-batch algorithms, we plot the validation log-likelihood vs. the total number of consumed samples. Results are shown in Figure 2. First, we observe that EM-based algorithms perform better than the gradient-based method Adam, suggesting that EM is a better parameter learning algorithm for PCs. Next, the proposed mini-batch EM algorithm outperforms both the full-batch EM algorithm and the mini-batch EM algorithm proposed by Peharz et al. [2020], indicating the empirical superiority of the proposed algorithm.

### v2

@@ Appendix C Global Parameter Renormalization of PCs
$\frac{\partial\log\tilde{{p}}_{\boldsymbol{\phi}^{\prime}}^{n}(\boldsymbol{x})}{\partial\log\tilde{{p}}_{\boldsymbol{\phi}^{\prime}}^{c}(\boldsymbol{x})}$ $=\frac{\tilde{{p}}_{\boldsymbol{\phi}^{\prime}}^{c}(\boldsymbol{x})}{\tilde{{p}}_{\boldsymbol{\phi}^{\prime}}^{n}(\boldsymbol{x})}\cdot\frac{\partial\tilde{{p}}_{\boldsymbol{\phi}^{\prime}}^{n}(\boldsymbol{x})}{\partial\tilde{{p}}_{\boldsymbol{\phi}^{\prime}}^{c}(\boldsymbol{x})},$
$=\frac{\tilde{{p}}_{\boldsymbol{\phi}^{\prime}}^{c}(\boldsymbol{x})}{\tilde{{p}}_{\boldsymbol{\phi}^{\prime}}^{n}(\boldsymbol{x})}\cdot\theta^{\prime}_{n,c},$
$=\frac{\tilde{{p}}_{\boldsymbol{\phi}}^{c}(\boldsymbol{x})/Z_{\boldsymbol{\phi}}(c)}{\tilde{{p}}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x})/Z_{\boldsymbol{\phi}}(n)}\cdot\theta^{\prime}_{n,c},$
$=\frac{\tilde{{p}}_{\boldsymbol{\phi}}^{c}(\boldsymbol{x})/\cancel{Z_{\boldsymbol{\phi}}(c)}}{\tilde{{p}}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x})/\cancel{Z_{\boldsymbol{\phi}}(n)}}\cdot\frac{\theta_{n,c}\cdot\cancel{Z_{\boldsymbol{\phi}}(c)}}{\cancel{Z_{\boldsymbol{\phi}}(n)}},$
$=\frac{\tilde{{p}}_{\boldsymbol{\phi}}^{c}(\boldsymbol{x})}{\tilde{{p}}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x})}\cdot\theta_{n,c},$
$=\frac{\tilde{{p}}_{\boldsymbol{\phi}}^{c}(\boldsymbol{x})}{\tilde{{p}}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x})}\cdot\frac{\partial\tilde{{p}}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x})}{\partial\tilde{{p}}_{\boldsymbol{\phi}}^{c}(\boldsymbol{x})},$
$=\frac{\partial\log\tilde{{p}}_{\boldsymbol{\phi}}^{c}(\boldsymbol{x})}{\partial\log\tilde{{p}}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x})}.$
Appendix D Additional Experimental Details
D.1 Details about the Datasets and the PC Models
ImageNet32 and The Corresponding PCs.
For ImageNet32, we partition every $32\times 32$ image (three color channels) into four $16\times 16$ patches and treat these as individual data samples. There are in total $16\times 16\times 3=768$ categorical variables in the PC.
We preprocess the data in color space with a lossy transformation, YCoCg, and its scaled, reversible variant, YCoCg-R, proposed by Malvar and Sullivan, (2003). Specifically, in YCoCg transformation, given a pixel with RGB values $(R,G,B)$, we first normalize them to the range $[0,1]$ by
@@ D.1 Details about the Datasets and the PC Models
where $y$, coco, and $cg$ are all in the range [1,1][-1,1]. Finally, we quantize the interval [1,1][-1,1] into $256$ bins uniformly and convert $y$, coco, and $cg$ to their quantized version $Y$, CoCo, and $Cg$, respectively. Note that $Y$, CoCo, and $Cg$ are all categorical variables with $256$ categories.
And the other transformation, YCoCg-R, maps 8-bit integer RGB values to YCoCg values without information loss. Similarly, the forward transformation is given by:
$co=r-b,\;tmp=b+co/2,\;cg=g-tmp,\;y=tmp+cg/2,$
where $y$, coco, and $cg$ are also in the range [1,1][-1,1]. The resulting integer channels are treated as categorical variables with $512$ categories.
We train several deep PC architectures on the ImageNet32 dataset Deng et al., (2009). These include Hidden Chow-Liu Trees (HCLT) (Liu and Van den Broeck,, 2021) with hidden size 512 and 1024, and Partitioned Data HCLTs (PDHCLTs) with 256 and 512 latents. [^9] The PDHCLT structure is described in Appendix E. The PDHCLT models are configured to partition the input data, which has a shape of (3, 16, 16). They use a maximum of 8 connections between product blocks. Please refer to
WikiText and The Corresponding PCs.
We also extend our empirical evaluation to language modeling using the WikiText-103 dataset by Merity et al., (2017). The raw text data is preprocessed into a format suitable for sequence modeling. Specifically, we firstly tokenize the entire corpus using the standard GPT-2 tokenizer by Radford et al., 2019b . All tokenized documents are then concatenated into a single continuous stream of tokens. Finally, this stream is partitioned into non-overlapping sequences of a fixed length of 128 tokens, with any remaining tokens at the end discarded to ensure uniformity across samples. On this preprocessed data, we trained Hidden Markov Models (HMMs), with 256, 512, and 1024 hidden states, and Monarch HMM with a size of 1024 (Zhang et al.,, 2025).
Biobank dataset and The Corresponding PCs.
For our bioinformatics experiments, we use genetic data sourced from the UK Biobank (UKBB) dataset (Sudlow et al.,, 2015). This specific version focuses on a single Linkage Disequilibrium (LD) block located on chromosome 6. The authors of this dataset remove SNPs with more than 1% missingness and those that deviate significantly from Hardy-Weinberg Equilibrium ($1\times 10^{-7}$ significance). Further, only individuals with no genetic relatedness to other individuals are considered. We have gotten the approval from the UK BioBank to access this dataset.
D.2 Details about the Optimizers
@@
The full-batch EM implementation follows prior work (e.g., Choi et al., (2021); Peharz et al., (2020)).
Anemone.
For notation simplicity, we define $\alpha=\eta/(\eta-1)$. Therefore, we can rewrite Equation 7 equivalently as
@@
which makes it more consistent with the baseline mini-batch EM algorithm.
We performed a preliminary hyperparameter search where we experimented with batch sizes including 512 and 16384. For the learning rate, we tested fixed values of $\alpha\in\{0.1,0.2,0.4,0.6\}$ and also employed a cosine decay schedule. The schedules included decreasing the rate from a base of $\alpha=0.4$ to a final rate of $\alpha=0.2$, and from $\alpha=0.8$ down to $\alpha=0.6$. We also set a momentum of $0.9$ in pratice. We select the final hyperparameter setting based on performance after the first 100 epochs.
@@
Following Loconte et al., (2025, 2024), we adopt the Adam optimizer (Kingma,, 2014). We selected hyperparameters using a similar search criterion as our EM experiments, testing learning rates of $\{1\times 10^{-2},3\times 10^{-3},5\times 10^{-3},1\times 10^{-3}\}$ and batch sizes of 512 and 1024. On the ImageNet32 YCC dataset, we found a learning rate of $1\times 10^{-2}$ performed the best, which aligns with the observations in Loconte et al., (2024). To ensure correctness, we first validated our implementation by reproducing prior results on the MNIST dataset, achieving a log-likelihood of -661.6 after 30 epochs.
D.3 Details about computing resources
We ran all the experiments included on NVIDIA A40s and NVIDIA GeForce RTX 4090.
D.4 Convergence experiments
We provide a similar experiment to Table 2, to demonstrate the consistency of faster convergence speed on various datasets, as shown in Table 5.
Table 5: Convergence speed (epochs) for HMM 256 on WikiText. The table reports epochs to reach specific $L$ thresholds, with $\Delta$ representing the difference from the best $L$ of -722 (Table 4). Lower is better. Bold marks the best result per column; $\infty$ indicates failure to reach the threshold in time.
Method $L$ $\geq$ -730 $L$ $\geq$ -724 $L$ $\geq$ -723
($\Delta\approx 7.8$) ($\Delta\approx 1.8$) ($\Delta\approx 0.8$)
Full EM 60 230 375
Adam $\infty$ $\infty$ $\infty$
Mini EM 45 $\infty$ $\infty$
Anemone 30 115 275
Appendix E The PDHCLT Structure
We adopt the PDHCLT structure implemented in the PyJuice (Liu et al.,, 2024) package (pyjuice.structures.PDHCLT). In the case of images with size $(3,16,16)$, we define a “split interval” to be $(3,4,4)$, which means that we partition the image into chunks of size $3\times 4\times 4$, resulting in $4\times 4$ chunks. For each chunk, we adopt the HCLT structure, and the PD structure is used to connect the different chunks.
For the BioBank dataset, the sequences have length $1167$, and we partition them into chunks of size $128$ (the last chunk has a smaller size). Again, HCLT is used to model intra-chunk dependencies and the PD structure is used to capture inter-chunk dependencies.